What monthly deposit must be made in order to accumulate 82,228 dollars in 12 years if a fund pays 8% per year, compounded monthly? Assume that deposits are made at the end of each month.
Answer #1

ANSWER:

I = 8% PER YEAR OR 8% / 12 PER MONTH

N = 12 YEARS OR 12 * 12 = 144 MONTHS

FV = $82,228

Monthly deposit = fv(a/f,i,n) = 82,228(a/f,8%/12,144) = 82,228 * 0.0042 = 341.89

so the monthly deposit is $341.89